A car takes off from rest and covers 25 m while accelerating at 2.0 m/s^2. How long does it take? What speed is obtained?
Answer by Alan3354(69443)

s = at^2/2
25 = 2t^2/2 = t^2
t = 5 seconds
-------------
v = at
v = 2*5 = 10 m/sec